# spacing=nonfrench

### v1

@@ IV.D  Properties (i)–(iii) are sufficient
Assume that a system with Hamiltonian $H$ is in a state $\rho$ whose support is contained in the direct sum of the eigenspaces associated with two energy eigenvalues $E_{0}$ and $E_{1}$, with $E_{0}<E_{1}$. Further assume that every eigenvector of $\rho$ with a nonzero eigenvalue admits a representation of the form in Eq. (12), so that the corresponding pure state evolves into a state with overlap $\delta$ after a time $\tau$ that saturates the Margolus–Levitin bound. Finally, let $\rho$ have a spectral decomposition of the form in Eq. (13), and assume that the eigenvectors in this decomposition evolve within mutually orthogonal two-dimensional subspaces. Under these conditions,
$\begin{split}\sqrt{\rho}\,U_{\tau}\sqrt{\rho}&=\sum_{j=1}^{r}\sum_{k=1}^{r}\sqrt{p_{j}p_{k}}\,|\psi_{j}\rangle\langle\psi_{j}|U_{\tau}|\psi_{k}\rangle\langle\psi_{k}|\\ &=\sum_{j=1}^{r}p_{j}|\psi_{j}\rangle\langle\psi_{j}|U_{\tau}|\psi_{j}\rangle\langle\psi_{j}|\\ &=\frac{1}{2}\big((1-z_{\delta})e^{-i\tau E_{0}}+(1+z_{\delta})e^{-i\tau E_{1}}\big)\rho.\end{split}$ (31)
@@ V  Discussion
The Margolus–Levitin quantum speed limit applies to both pure and mixed states, and in this work we have identified the precise circumstances under which it can be saturated. Specifically, if a system evolves unitarily from an initial state $\rho$ to a state with fidelity $\delta$ in a time $\tau$ that attains the Margolus–Levitin bound, then $\rho$ must be a convex combination of pure states that populate the same two energy eigenvalues, have identical expected energy, and individually saturate the bound with overlap $\delta$ at time $\tau$. Moreover, the supports of these pure states must evolve within mutually orthogonal subspaces.
@@
This modification naturally raises the question of whether Uhlmann–Jozsa fidelity is the most appropriate extension of the pure-state overlap to mixed states in this setting. While fidelity is a standard and operationally meaningful choice, it is by no means unique; several alternative generalizations of overlap exist and may lead to distinct or potentially sharper speed limits. Investigating such alternatives will be the subject of future work.
@@ V.A  Quantum bits
$H=E_{0}|E_{0}\rangle\langle E_{0}|+E_{1}|E_{1}\rangle\langle E_{1}|,\quad E_{0}<E_{1}.$ (35)
@@
$x$ $=\langle E_{1}|\rho|E_{0}\rangle+\langle E_{0}|\rho|E_{1}\rangle,$ (36a)
$y$ $=i\!\left(\langle E_{1}|\rho|E_{0}\rangle-\langle E_{0}|\rho|E_{1}\rangle\right),$ (36b)
$z$ $=\langle E_{1}|\rho|E_{1}\rangle-\langle E_{0}|\rho|E_{0}\rangle.$ (36c)
@@
The dynamics generated by $H$ cause the Bloch vector to rotate about the $z$-axis with constant inclination and angular speed $\omega=E_{1}-E_{0}$. Let $\bm{r}_{\tau}$ denote the Bloch vector of the state at time $\tau$. During the evolution, the orthogonal projection of the Bloch vector onto the $xy$-plane rotates through an angle
@@
$\tau=\frac{\theta}{\omega}\geq\frac{1}{E_{1}-E_{0}}\,\arccos\!\bigg(\frac{\bm{r}\cdot\bm{r}_{\tau}-z^{2}}{2\wp-1-z^{2}}\bigg).$ (40)
@@
Assume that the fidelity between $\rho$ and $\rho_{\tau}$ is $\delta$. Hübner’s formula for the fidelity between qubit states [30] gives
$\delta=\operatorname{tr}(\rho\rho_{\tau})+2\sqrt{\det\rho\,\det\rho_{\tau}}\\ =\frac{1}{2}\!\left(3+\bm{r}\!\cdot\!\bm{r}_{\tau}-2\wp\right).$ (41)
@@
and the bound (45) is saturated.
By adjusting the phases of the energy eigenvectors $|E_{0}\rangle$ and $|E_{1}\rangle$, if necessary, we may take the initial Bloch vector to have $x$-component $(2\wp-1-z_{\delta,\wp}^{2})^{1/2}$ and $y$-component 0. With this choice of phases, the state assumes the form
$\begin{split}\rho=&\frac{1}{2}(1-z_{\delta,\wp})|E_{0}\rangle\langle E_{0}|\\ &\,+\frac{1}{2}\sqrt{2\wp-1-z_{\delta,\wp}^{2}}\,\big(|E_{0}\rangle\langle E_{1}|+|E_{1}\rangle\langle E_{0}|\big)\\ &\,+\frac{1}{2}(1+z_{\delta,\wp})|E_{1}\rangle\langle E_{1}|.\end{split}$ (47)
In the pure-state limit, $\rho=|\psi\rangle\langle\psi|$, where
$|\psi\rangle=\sqrt{\frac{1-z_{\delta,1}}{2}}|E_{0}\rangle+\sqrt{\frac{1+z_{\delta,1}}{2}}|E_{1}\rangle.$ (48)
@@ V.B  The dual Margolus–Levitin speed limit
If the inequality in Eq. (50) is saturated, then the support of the initial state is necessarily contained within the eigenspaces of $H$ corresponding to the energies $E_{0}$ and $E_{m}$. Moreover, each eigenvector of $\rho$ with a nonzero eigenvalue admits a representation of the form
@@ Appendix D Proof that Equation (22) implies Equation (23)
Let $P_{0}$ and $P_{1}$ be the orthogonal projections onto the eigenspaces $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ of $H$ corresponding to eigenvalues $E_{0}$ and $E_{1}$. Since the support of $\rho$, and thus of $\sqrt{\rho}$, is contained in $\mathcal{E}_{0}\oplus\mathcal{E}_{1}$, we have
$\rho=\sqrt{\rho}P_{0}\sqrt{\rho}+\sqrt{\rho}P_{1}\sqrt{\rho},$ (60a)
$\sqrt{\rho}\,U_{\tau}\sqrt{\rho}=e^{-i\tau E_{0}}\sqrt{\rho}P_{0}\sqrt{\rho}+e^{-i\tau E_{1}}\sqrt{\rho}P_{1}\sqrt{\rho}.$ (60b)
@@
$\sqrt{\rho}\,U_{\tau}\sqrt{\rho},\sqrt{\rho}\,U_{\tau}^{\dagger}\sqrt{\rho}]=2i\sin(\tau(E_{1}-E_{0}))\,[\sqrt{\rho}P_{0}\sqrt{\rho},\sqrt{\rho}P_{1}\sqrt{\rho}\,].$ (61)
Assume that $e^{-i\tau E_{0}}\neq-e^{-i\tau E_{1}}$, so that $\sin(\tau(E_{1}-E_{0}))\neq 0$. (The case $e^{-i\tau E_{0}}=e^{-i\tau E_{1}}$ is trivial, since $U_{\tau}$ then leaves $\rho$ invariant.) Then $\sqrt{\rho}P_{0}\sqrt{\rho}$ and $\sqrt{\rho}P_{1}\sqrt{\rho}$ must commute. Consequently, there exists an orthonormal basis $|\phi_{1}\rangle,|\phi_{2}\rangle,\dots,|\phi_{r}\rangle$ for the support of $\rho$, consisting of simultaneous eigenvectors of $\sqrt{\rho}P_{0}\sqrt{\rho}$ and $\sqrt{\rho}P_{1}\sqrt{\rho}$, with
@@
$\sqrt{\rho}P_{1}\sqrt{\rho}|\phi_{j}\rangle$ $=q_{j}^{1}\,|\phi_{j}\rangle.$ (62b)
The eigenvalues $q_{j}^{0}$ and $q_{j}^{1}$ are nonnegative, since both operators are positive semidefinite. Moreover, $|\phi_{j}\rangle$ is an eigenvector of $\rho$ with eigenvalue $p_{j}=q_{j}^{0}+q_{j}^{1}$, which is strictly positive since $|\phi_{j}\rangle$ belongs to the support of $\rho$. Define $s_{j}=q_{j}^{0}/p_{j}$. Then $0\leq s_{j}\leq 1$, $q_{j}^{0}=p_{j}s_{j}$, and $q_{j}^{1}=p_{j}(1-s_{j})$. Introduce the function
$q(s)=se^{-i\tau E_{0}}+(1-s)e^{-i\tau E_{1}},$ (63)
which traces the segment between $e^{-i\tau E_{0}}$ and $e^{-i\tau E_{1}}$ in the complex plane as $s$ runs over the interval $[0,1]$. Then $|\phi_{j}\rangle$ is an eigenvector of $\sqrt{\rho}\,U_{\tau}\,\sqrt{\rho}$ with eigenvalue $p_{j}q(s_{j})$:
$\sqrt{\rho}\,U_{\tau}\sqrt{\rho}\,|\phi_{j}\rangle=(e^{-i\tau E_{0}}\sqrt{\rho}P_{0}\sqrt{\rho}+e^{-i\tau E_{1}}\sqrt{\rho}P_{1}\sqrt{\rho})|\phi_{j}\rangle=p_{j}q(s_{j})|\phi_{j}\rangle.$ (64)
The condition $\sqrt{\rho}U_{\tau}\sqrt{\rho}=e^{i\theta}|\sqrt{\rho}U_{\tau}\sqrt{\rho}|$ requires all these eigenvalues to share the same phase factor $e^{i\theta}$. Thus, the complex numbers $q(s_{j})$ must lie on a common ray from the origin. At the same time, each lies on the line segment connecting $e^{-i\tau E_{0}}$ and $e^{-i\tau E_{1}}$, which does not pass through the origin under our assumption $e^{-i\tau E_{0}}\neq-e^{-i\tau E_{1}}$. Hence, the ray and the segment intersect at a unique point $q$, implying that $q(s_{j})=q$ for all $j$. Then
@@
It remains to analyze the special case $e^{-i\tau E_{0}}=-e^{-i\tau E_{1}}$. Since the purified system described at the beginning of Sec. IV.3 saturates the Margolus–Levitin quantum speed limit, the purification $|w\rangle$ can be written as
$|w\rangle=\sqrt{\frac{1-z_{\delta}}{2}}|E_{0}\rangle+\sqrt{\frac{1+z_{\delta}}{2}}|E_{1}\rangle,$ (67)
where $|E_{0}\rangle$ and $|E_{1}\rangle$ are normalized vectors in $\mathcal{E}_{0}\otimes\mathcal{H}$ and $\mathcal{E}_{1}\otimes\mathcal{H}$, respectively. Then
$\sqrt{\delta}=|\langle w|w_{\tau}\rangle|=\frac{1}{2}\big|(1-z_{\delta})e^{-i\tau E_{0}}+(1+z_{\delta})e^{-i\tau E_{1}}\big|=|-e^{-i\tau E_{0}}z_{\delta}|=-z_{\delta}.$ (68)

### v2

@@ IV.D  Properties (i)–(iii) are sufficient
Assume that a system with Hamiltonian $H$ is in a state $\rho$ whose support is contained in the sum of the eigenspaces associated with two energy eigenvalues $E_{0}$ and $E_{m}$, with $E_{0}<E_{m}$. Further assume that every eigenvector of $\rho$ with a nonzero eigenvalue admits a representation of the form in Eq. (12), so that the corresponding pure state evolves into a state with overlap $\delta$ after a time $\tau$ that saturates the Margolus–Levitin bound. Finally, let $\rho$ have a spectral decomposition of the form in Eq. (13), and assume that the eigenvectors in this decomposition evolve within mutually orthogonal subspaces. Under these conditions,
$\begin{split}\sqrt{\rho}\,U_{\tau}\sqrt{\rho}&=\sum_{j=1}^{r}\sum_{k=1}^{r}\sqrt{p_{j}p_{k}}\,|\psi_{j}\rangle\langle\psi_{j}|U_{\tau}|\psi_{k}\rangle\langle\psi_{k}|\\ &=\sum_{j=1}^{r}p_{j}|\psi_{j}\rangle\langle\psi_{j}|U_{\tau}|\psi_{j}\rangle\langle\psi_{j}|\\ &=\frac{1}{2}\big((1-z_{\delta})e^{-i\tau E_{0}}+(1+z_{\delta})e^{-i\tau E_{m}}\big)\rho.\end{split}$ (31)
@@ V  Discussion
The Margolus–Levitin quantum speed limit (3) applies to both pure and mixed states, and in this work we have identified the precise circumstances under which it can be saturated. Specifically, if a system evolves unitarily from an initial state $\rho$ to a state with fidelity $\delta$ in a time $\tau$ that attains the Margolus–Levitin bound, then $\rho$ must be a convex combination of pure states that populate the same two energy eigenvalues, have identical expected energy, and individually saturate the bound with overlap $\delta$ at time $\tau$. Moreover, the supports of these pure states must evolve within mutually orthogonal subspaces.
@@
This modification naturally raises the question of whether Uhlmann–Jozsa fidelity is the most appropriate extension of the pure-state overlap to mixed states in this setting. While fidelity is a standard and operationally meaningful choice, it is by no means unique; several alternative generalizations of overlap exist and may yield distinct or sharper speed limits. Investigating such alternatives will be the subject of future work.
@@ V.A  Quantum bits
$H=E_{0}|E_{0}\rangle\langle E_{0}|+E_{m}|E_{m}\rangle\langle E_{m}|,\quad E_{0}<E_{m}.$ (35)
@@
$x$ $=\langle E_{m}|\rho|E_{0}\rangle+\langle E_{0}|\rho|E_{m}\rangle,$ (36a)
$y$ $=i\!\left(\langle E_{m}|\rho|E_{0}\rangle-\langle E_{0}|\rho|E_{m}\rangle\right),$ (36b)
$z$ $=\langle E_{m}|\rho|E_{m}\rangle-\langle E_{0}|\rho|E_{0}\rangle.$ (36c)
@@
The dynamics generated by $H$ cause the Bloch vector to rotate about the $z$-axis with constant inclination and angular speed $\omega=E_{m}-E_{0}$. Let $\bm{r}_{\tau}$ denote the Bloch vector of the state at time $\tau$. During the evolution, the orthogonal projection of the Bloch vector onto the $xy$-plane rotates through an angle
@@
$\tau=\frac{\theta}{\omega}\geq\frac{1}{E_{m}-E_{0}}\,\arccos\!\bigg(\frac{\bm{r}\cdot\bm{r}_{\tau}-z^{2}}{2\wp-1-z^{2}}\bigg).$ (40)
@@
Assume that the fidelity between $\rho$ and $\rho_{\tau}$ is $\delta$. Hübner’s formula [30] for the fidelity between qubit states gives
$\delta=\operatorname{tr}(\rho\rho_{\tau})+2\sqrt{\det\rho\,\det\rho_{\tau}}=\frac{1}{2}(3+\bm{r}\cdot\bm{r}_{\tau}-2\wp).$ (41)
@@
and the bound in Eq. (45) is saturated.
By adjusting the phases of the energy eigenvectors $|E_{0}\rangle$ and $|E_{m}\rangle$, if necessary, we may take the initial Bloch vector to have $x$-component $(2\wp-1-z_{\delta,\wp}^{2})^{1/2}$ and $y$-component 0. With this choice of phases, the state assumes the form
$\begin{split}\rho=&\frac{1}{2}(1-z_{\delta,\wp})|E_{0}\rangle\langle E_{0}|\\ &\,+\frac{1}{2}\sqrt{2\wp-1-z_{\delta,\wp}^{2}}\,\big(|E_{0}\rangle\langle E_{m}|+|E_{m}\rangle\langle E_{0}|\big)\\ &\,+\frac{1}{2}(1+z_{\delta,\wp})|E_{m}\rangle\langle E_{m}|.\end{split}$ (47)
In the pure-state limit we have that $\rho=|\psi\rangle\langle\psi|$, where
$|\psi\rangle=\sqrt{\frac{1-z_{\delta,1}}{2}}|E_{0}\rangle+\sqrt{\frac{1+z_{\delta,1}}{2}}|E_{m}\rangle.$ (48)
@@
Remark 3.
The bound in Eq. (40) is reminiscent of the operational quantum speed limit for systems with time-independent Hamiltonians introduced in Ref. [31] and further studied in Ref. [32]. In fact, Eq. (40) may be regarded as a refinement of the operational quantum speed limit obtained by explicitly accounting for energy conservation. A more detailed analysis of how conserved quantities can be systematically incorporated into the operational framework to obtain progressively stronger lower bounds on the evolution time will be presented elsewhere.
@@ V.B  The dual Margolus–Levitin speed limit
If the inequality in Eq. (50) is saturated, then the support of the initial state is necessarily contained within the eigenspaces of $H$ corresponding to the energies $E_{0}$ and $E_{m}$. Moreover, each eigenvector $|\psi\rangle$ of $\rho$ with a nonzero eigenvalue admits a representation of the form
@@
Remark 4.
The sign difference between Eqs. (12) and (52) implies that the families of states saturating the Margolus–Levitin quantum speed limit and its dual coincide if and only if $z_{\delta}=0$, which occurs only for $\delta=0$. Otherwise, the two families are disjoint. Consequently, the two bounds can be saturated simultaneously only when the fidelity between the initial and final states is zero, in which case saturation of one bound implies saturation of the other.
@@ Appendix D Proof that Equation (22) implies Equation (23)
Let $P_{0}$ and $P_{m}$ be the orthogonal projections onto the eigenspaces $\mathcal{E}_{0}$ and $\mathcal{E}_{m}$ of $H$ corresponding to eigenvalues $E_{0}$ and $E_{m}$. Since the support of $\rho$, and thus of $\sqrt{\rho}$, is contained in $\mathcal{E}_{0}\oplus\mathcal{E}_{m}$, we have
$\rho=\sqrt{\rho}P_{0}\sqrt{\rho}+\sqrt{\rho}P_{m}\sqrt{\rho},$ (60a)
$\sqrt{\rho}\,U_{\tau}\sqrt{\rho}=e^{-i\tau E_{0}}\sqrt{\rho}P_{0}\sqrt{\rho}+e^{-i\tau E_{m}}\sqrt{\rho}P_{m}\sqrt{\rho}.$ (60b)
@@
$\sqrt{\rho}\,U_{\tau}\sqrt{\rho},\sqrt{\rho}\,U_{\tau}^{\dagger}\sqrt{\rho}]=2i\sin(\tau(E_{m}-E_{0}))\,[\sqrt{\rho}P_{0}\sqrt{\rho},\sqrt{\rho}P_{m}\sqrt{\rho}\,].$ (61)
Assume that $e^{-i\tau E_{0}}\neq-e^{-i\tau E_{m}}$, so that $\sin(\tau(E_{m}-E_{0}))\neq 0$. (The case $e^{-i\tau E_{0}}=e^{-i\tau E_{m}}$ is trivial since $U_{\tau}$ then leaves $\rho$ invariant.) Then $\sqrt{\rho}P_{0}\sqrt{\rho}$ and $\sqrt{\rho}P_{m}\sqrt{\rho}$ must commute. Consequently, there exists an orthonormal basis $|\phi_{1}\rangle,|\phi_{2}\rangle,\dots,|\phi_{r}\rangle$ for the support of $\rho$, consisting of simultaneous eigenvectors of $\sqrt{\rho}P_{0}\sqrt{\rho}$ and $\sqrt{\rho}P_{m}\sqrt{\rho}$, with
@@
$\sqrt{\rho}P_{m}\sqrt{\rho}|\phi_{j}\rangle$ $=q_{j}^{m}\,|\phi_{j}\rangle.$ (62b)
The eigenvalues $q_{j}^{0}$ and $q_{j}^{m}$ are nonnegative since both operators are positive semidefinite. Moreover, $|\phi_{j}\rangle$ is an eigenvector of $\rho$ with eigenvalue $p_{j}=q_{j}^{0}+q_{j}^{m}$, which is strictly positive since $|\phi_{j}\rangle$ belongs to the support of $\rho$. Define $s_{j}=q_{j}^{0}/p_{j}$. Then $0\leq s_{j}\leq 1$, $q_{j}^{0}=p_{j}s_{j}$, and $q_{j}^{m}=p_{j}(1-s_{j})$. Introduce the function
$q(s)=se^{-i\tau E_{0}}+(1-s)e^{-i\tau E_{m}},$ (63)
which traces the segment between $e^{-i\tau E_{0}}$ and $e^{-i\tau E_{m}}$ in the complex plane as $s$ runs over the interval $[0,1]$. Then $|\phi_{j}\rangle$ is an eigenvector of $\sqrt{\rho}\,U_{\tau}\,\sqrt{\rho}$ with eigenvalue $p_{j}q(s_{j})$:
$\sqrt{\rho}\,U_{\tau}\sqrt{\rho}\,|\phi_{j}\rangle=(e^{-i\tau E_{0}}\sqrt{\rho}P_{0}\sqrt{\rho}+e^{-i\tau E_{m}}\sqrt{\rho}P_{m}\sqrt{\rho})|\phi_{j}\rangle=p_{j}q(s_{j})|\phi_{j}\rangle.$ (64)
The condition $\sqrt{\rho}U_{\tau}\sqrt{\rho}=e^{i\theta}|\sqrt{\rho}U_{\tau}\sqrt{\rho}|$ requires all these eigenvalues to share the same phase factor $e^{i\theta}$. Thus, the complex numbers $q(s_{j})$ must lie on a common ray from the origin. At the same time, each lies on the line segment connecting $e^{-i\tau E_{0}}$ and $e^{-i\tau E_{m}}$, which does not pass through the origin under our assumption $e^{-i\tau E_{0}}\neq-e^{-i\tau E_{m}}$. Hence, the ray and the segment intersect at a unique point $q$, implying that $q(s_{j})=q$ for all $j$. Then
@@
It remains to analyze the special case $e^{-i\tau E_{0}}=-e^{-i\tau E_{m}}$. Since the purified system described at the beginning of Sec. IV.3 saturates the Margolus–Levitin quantum speed limit, the purification $|w\rangle$ can be written as
$|w\rangle=\sqrt{\frac{1-z_{\delta}}{2}}|E_{0}\rangle+\sqrt{\frac{1+z_{\delta}}{2}}|E_{m}\rangle,$ (67)
where $|E_{0}\rangle$ and $|E_{m}\rangle$ are normalized vectors in $\mathcal{E}_{0}\otimes\mathcal{H}$ and $\mathcal{E}_{m}\otimes\mathcal{H}$, respectively. Then
$\sqrt{\delta}=|\langle w|w_{\tau}\rangle|=\frac{1}{2}\big|(1-z_{\delta})e^{-i\tau E_{0}}+(1+z_{\delta})e^{-i\tau E_{m}}\big|=|-e^{-i\tau E_{0}}z_{\delta}|=-z_{\delta}.$ (68)